The Basics of Magic in Lords 32

Ryan Powers

September 15, 2001

1 Introduction

Magic and the supernatural exist in this campaign. However, it will be rare at the beginning because magic is just beginning to return to this world. This means that the world is not crawling with magical beast, artifacts, or even magicians themselves. However, the level of magic is rising, and this will cause some or all of these things to occur at some point. This means that every nation, religion, order, and secret empire will have to forge its own path to the supernatural, or ignore it completely if so desired.

2 New Terms and Concepts

Global Magic Level A measure of the magical energy level the world has reached. Increasing over time.

Regional Magic Level Local fluctuations on the World Magic Level.

Threshold Magic Level The minimum Regional Magic Level in the target region(s) required to perform a given magical action.

Magical Tech Level This is analogous to the normal Tech Level of the nation.

Magical Force Points A measure of the magical ability of the citizenry as a whole. Used to assist in large scale magical operations.

Magical Operations Large scale magical feats, ie raising armies, destroying whole regions, increasing production, etc.

Magic Rating A Leader statistic concerning their proficiency in magic.

Magical Leader Actions New leader actions using the Magic Rating of the leader.

Religion and Magic The religions of the world will have to deal with magic now.

3 Global, Regional and Threshold Magic Levels

The amount of magical energy available in the world is not a constant. It is growing. This is represented by the Global Magic Level. This level will start at one and increase from there, the approximate value of this level will be reported in the newsfax.

It is possible that their may be regional fluctuations to this level. These regional fluctuations are represented by the Regional Magic Level.

Many magical actions have a Threshold Magic Level, which is the minimum Regional Magic Level that must be reached before this ability is possible in the region. Regional Magic Levels above this value will provide a bonus to the action.

4 Magical Tech Level

Each Nation will have a magical Tech Level. The tech level will determine what operations the nation can learn to use. This can be increases with magical research. See the research document for more details.

5 Magical Force Points

Nations are able to produce Magical Force Points, MFP, which can be used to invest in various aspects of magic or used to aid in magical operations much like GP can be used to benefit Espionage Operations. For more details regarding magical operations see that section.

5.1 Producing MFP

MFP are produced at a base rate equal to the GPv of a Region, or /emphtwice the GPv of a city. This is modified by the normal tax rate modifiers regarding control status. Here are the relevant formulas for production.

T otalMF P = (RegionalMF P + CityMF P) × T axRate × RT F

RT F = 1 + \frac{(Religion′sT olerance − 6) × RS}{100}

RegionalMF P = ∑ RegionalGP v × StatusMultiplier

CityMF P = ∑ CityGP v × StatusMultiplier

TotalMFP The Total MFP production of the nation. RegionalMFP The total regional production of MFP. CityMFP The total city production of MFP. TaxRate The nations Tax Rate, modified by things like famine, census,

etc. RTF Religious Tolerance Factor, the variation based on the religions ac-

ceptance of magic. RS Religious Strength RegionalGPv The GPv of a region. CityGPv the GPv of a city. StatusMultiplier The status multiplier found in the base rules on table

2–13.

5.2 Research with MFP

In addition to aiding investing in ratings and using magical operations, MFP can be used in more focused research. This can be used to learn an ability that you have knowledge of. It can be used to learn of an ability. It can also be used to research an entirely new ability, let me know about this and we’ll work something out. This way different nations will have entirely different sets of magical possibilities.

6 Magical Operations

Magical operations work much like espionage, or religious actions. These operations rely on the nations Magical Operations Capacity, and the Magical Operations Bonus. These actions tend to represent large scale and/or long term magical activities. Some possibilities include:

• increasing production of Agro, NFP, MFP, and/or GP.

• rasing an army of magical creatures

• providing large scale assistance to an existing army by increasing offensive or defensive capabilities.

• changing of terrain types

These Operations will have a minimum MOB point cost to use the ability. MFP may be added to these operations just like GP to Espionage and Religious Actions. For the Operations than can have a duration spanning multiple turns, they will have a support cost (usually in MFP).

6.1 Learning Operations

Knowledge of magical operations may be given to another nation, this gives them a very large bonus in learning it, but does not confer automatic success. Their existing views and perspectives on how magic works may interfere.

For more information on learning new operations, check out the research document.

6.2 Magical Operations Capacity

Every nation has a Magical Operations Capacity, or MOC. This is the number of Operations that can be performed in a turn. Each Operation costs exactly one point of MOC. Much like the espionage and religious capacities. This value can be invested in just like the espionage and religious Capacity bonuses. MFP can also be invested in this in a 1MFP=1NFP=2GP ratio.

6.3 Magical Operations Bonus

Each Nation also has a Magical Operations Bonus, or MOB. This is a measure of the quality of the nations ability to perform magical operations. These bonus points allow larger operations to be undertaken, or allow a given operation to be undertaken with a greater chance of success, or greater effect.

Each Operation has a minimum number of Bonus Points needed in order to succeed. Points may be added above the minimum in order to increase the chances of success. Certain Operations may have other uses for MOB read the descriptions to find out.

This value can be invested in just like MOC points including the ability to invest MFP points in addition to NFP and GP.

6.4 Maintenance Costs

MOC and MOB points need to be maintained. The rate is the same as for religious operations. This cost is in GP.

MaintenaceCost = (MOC + MOB)^{2}

6.5 Magical Force Points in Operations

Magical Force Points have several uses with regard to magical operations. They can be used to add bonuses to the chance of success much the same as GP can be added to religious and espionage actions. They can also be used to sustain an existing operation over multiple turns. If this is possible for a given operation, the costs and effects will be described in the operation description.

7 Magic and Leaders

7.1 The Magic Rating

Each leader has a fifth stat, their magic rating. This is a measure of the leaders skill in manipulating magical energies. It is on the 1-B scale just like military ratings.

This rating is fixed, but it may be possible to pool the abilities of more than one leader to use a higher rating. In Order to do this a Magical Tech Level of at least two must be reached.

7.2 Magical Leader Actions

There are new Leader Actions which rely on the Magic Rating of the leader. Some of these will be known to everybody, the rest will need to be researched.

The knowledge of these abilities will be tradeable just like the Magical Operations.

7.3 New Leaders

A Wizard leader can be taken in place of a lieutenant. This leader will have a minimum Magic Rating based on the Magical Tech Level of the Nation.

8 Religion and Magic

Each Religion will have a magical tolerance rating of from 1 to 10. The lower this value is, the more opposed to magic that religions beliefs are. This can be changed through various actions.

8.1 MFP production and Religion

A high tolerance will produce a bonus to MFP production while a low tolerance will produce a penalty. This will be modified based on the Religious Strength of the nation. It ranges from +50% to -50%.

8.2 Magical Resistance and Religion

A low tolerance for magic will give a religions followers a certain level of resistance to some magical effects. While a high tolerance will provide a penalty to resisting such effects. Again the nation’s Religious Strength will affect the scale of this bonus or penalty.